# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ IV-B Survey on BD-RIS Optimization and Performance Analysis
Consider a general BD-RIS-aided multi-user MIMO system consisting of an $N$-antenna transmitter, an $M$-element BD-RIS, and $K$ multi-antenna users, each of which has $N_{k}$ antennas, $\forall k\in\mathcal{K}=\{1,\ldots,K\}$. Define the wireless channels from the transmitter to user $k$ as $\mathbf{H}_{{RT},k}\in\mathbb{C}^{N_{k}\times N}$, from the transmitter to BD-RIS as $\mathbf{H}_{{IT}}\in\mathbb{C}^{M\times N}$, and from the BD-RIS to user $k$ as $\mathbf{H}_{{RI},k}\in\mathbb{C}^{N_{k}\times M}$. According to [35], assuming perfect matching, no mutual coupling, no specular reflection, and unilateral approximation at all devices, the overall channel $\mathbf{H}_{k}(\mathbf{\Theta})\in\mathbb{C}^{N_{k}\times N}$ from the transmitter to user $k$, which is a function of the scattering matrix $\mathbf{\Theta}$ of BD-RIS, writes as
@@
where $F(\mathbf{W},\{\mathbf{H}_{k}(\mathbf{\Theta})\}_{\forall k})$ denotes a general utility function varying according to performance metrics, $\mathcal{Y}$ denotes the constraint of $\mathbf{Y}_{I}$ varying according to architectures, and $\mathcal{W}$ denotes the constraint of the precoder matrix varying according to budget requirements. Note that (62) can be used to optimize all reciprocal architectures in Section III-C where $\mathbf{Y}_{I}$ is explicitly defined and constrained. Alternatively, the optimization problem can be formulated as the following scattering matrix based form
@@
where $\mathcal{T}$ denotes the constraints of $\mathbf{\Theta}$ and varying according to architectures. Note that (63) can be used to optimize single-, group-, fully-connected architectures in Section III-C and non-reciprocal architectures in Section III-D.
@@
For reciprocal BD-RIS, with the form of (63) and the constraint of dynamically connected architectures, [44] has proposed an offline grouping strategy design adapting to the static CSI.
For non-reciprocal BD-RIS, the following beamforming design and performance analysis studies have been conducted.
  • With the form of (63) and a non-reciprocal BD-RIS architecture having non-diagonal scattering matrices being constrained by the set
@@
    closed-form solutions can be derived. Focusing primarily on SISO systems with the utility function $F_{\mathsf{SISO}}(h(\mathbf{\Theta}))$, [45] has derived the performance upper-bound and the closed-form solution for BD-RIS. The beamforming design is later extended to MISO systems with the utility function $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$.
  • With the form of (63), [46] has further proposed a coordinated non-reciprocal group-connected architecture with scattering matrices having more than $M$ non-zero entries, and derived the optimal grouping strategy which could maximize $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$. This architecture is further used to support the multi-sector mode and maximize $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$ related to specific users [58].
@@ IV-B.2 Capacity Maximization for MIMO
With the form of (63), the following beamforming design studies have been conducted.
  • Focusing on reciprocal and fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$ ($G=1$) with
@@
    a common strategy is to iteratively design BD-RIS and the covariance matrix [59]. Specifically, the BD-RIS design can generally follow the idea of decoupling $\mathbf{\Theta}$ into $\mathbf{\Theta}=\mathbf{\Psi}^{\mathsf{H}}\mathbf{\Psi}$ with $\mathbf{\Psi}$ being unitary. For more specific LoS MIMO channels (both $\mathbf{H}_{RI}$ and $\mathbf{H}_{IT}$ have rank 1), the closed-form solution of $\mathbf{\Theta}$ can be obtained by exploiting the rank-1 property of channels $\mathbf{H}_{RI}$ and $\mathbf{H}_{IT}$ [60].
  • Focusing on non-reciprocal BD-RIS architecture constrained by the set
@@
    the problem has been directly solved by the Manifold method [50]. In addition, matrix decomposition based solutions have been proposed in closed form considering either near-field [61] or far-field [62] scenarios, which are proved to be globally optimal.
@@ IV-B.3 Sum-Rate Maximization for Multi-User MISO
This is again a multi-variable optimization problem, which can be solved either by separately designing BD-RIS and transmit precoder, or by block coordinate descent (BCD) methods [63].
For BD-RIS having reciprocal architectures, the following beamforming design and performance analysis studies have been conducted.
  • With the form of (63) and group/fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$, a two-stage BD-RIS and precoder design has been proposed in [54] with low computational complexity. Specifically, the optimization of BD-RIS matrix is based on the symmetric unitary projection as detailed in Section IV-A. Following the idea of symmetric unitary projection, a joint precoder and BD-RIS design has been further proposed in [64].
  • With the form of (62) and group/fully-connected architectures constrained by $\mathcal{Y}_{\mathsf{group-conn}}$, a quasi-Newton based solution together with a heuristic user scheduling scheme has been proposed in [65].
  • With the form of (63) and multi-sector mode BD-RIS constrained by the set
@@
    where $\varTheta_{l,1,n}$, $\forall l\in\mathcal{L}$, $\forall n=1,\ldots,\frac{M}{L}$ are extracted from $\mathbf{\Theta}$ according to Example 2 of Section III-E, an iteratively closed-form solution can be obtained by deriving the Karush–Kuhn–Tucker conditions [38]. In addition, [66] has derived the closed-form expression of the achievable sum-rate aided by multi-sector BD-RIS, with a further extension to other performance metrics, such as energy efficiency, error probability, outage probability [67].
For BD-RIS having non-reciprocal architectures, the following beamforming design studies have been conducted.
  • For BD-RIS constrained by the set $\mathcal{T}_{\mathsf{non-recip}}$, [68] and [69] have proposed learning-based methods which support larger-dimensional optimizations.
  • For BD-RIS with hybrid mode and proper non-reciprocal architectures, the constraint $\mathbf{\Theta}_{\mathsf{r}}=\mathbf{\Theta}_{\mathsf{r}}^{\mathsf{T}}$ from Example 1 of Section III-E can be dropped such that $\mathbf{\Theta}$ is constrained by the set
@@
    where $\mathbf{\Theta}_{\mathsf{r},g}$ and $\mathbf{\Theta}_{\mathsf{t},g}$, $\forall g\in\mathcal{G}$ are extracted from $\mathbf{\Theta}$ according to Example 1 of Section III-E. It should be noted here that the constraint of hybrid BD-RIS essentially constructs a Stiefel Manifold [49]. Therefore, an iterative searching-based solution can be obtained by using the Manifold theory [36].
  • For BD-RIS with hybrid mode and dynamically connected architectures adapting to the instantaneous CSI, [43] has proposed a heuristic grouping strategy design.
@@ IV-B.4 Energy Efficiency Maximization and Power Minimization for Multi-User MISO
For BD-RIS having reciprocal architectures, the following beamforming design and performance analysis studies have been conducted.
  • With the form of (63) and group/fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$, [56] has proposed an iterative framework by decoupling the unitary and symmetric constraints of $\mathbf{\Theta}$ with a newly introduced auxiliary variable $\mathbf{\Phi}=\mathbf{\Theta}$, such that $\mathbf{\Theta}$ is only subject to the unitary constraint and $\mathbf{\Phi}$ is only subject to the symmetric constraint, or vise versa. The proposed framework is applicable to sum-rate maximization, energy-efficiency maximization, and power minimization problems.
  • With the form of (63) and STARS, which is essentially a special case of multi-sector BD-RIS constrained by $\mathcal{T}_{\mathsf{multi-sec}}$ when $L=2$ and also a special case of hybrid BD-RIS constrained by $\mathcal{T}_{\mathsf{hyb}}$ when $\bar{M}=2$, [70] has proposed a meta-learning approach for joint optimization of BD-RIS, resource allocation, and antenna selection.
  • With the form of (62), [57] has recently proposed a universal framework applicable to arbitrary constraints of BD-RIS admittance matrix, the aforementioned three utility functions, and sum-rate maximization for more general multi-user MIMO scenarios. The main idea is to introduce auxiliary variables $\mathbf{u}_{k}=\mathbf{h}_{RI,k}\mathbf{\Theta}$, $\forall k\in\mathcal{K}$, such that the matrix inverse in (62) can be eliminated by transferring it to bilinear constrains to facilitate the optimization. The applicable BD-RIS constraints include, but are not limited to, $\mathcal{Y}_{\mathsf{group-conn}}$ defined as
    $\mathcal{Y}_{\mathsf{group-conn}}=\{\mathbf{Y}_{I}=\mathsf{blkdiag}(\mathbf{Y}_{I,1},\ldots,\mathbf{Y}_{I,G})~{}|$ (78)
    $~{}~{}\Re\{\mathbf{Y}_{I,g}\}=\mathbf{0}_{\bar{M}\times\bar{M}},\mathbf{Y}_{I,g}=\mathbf{Y}_{I,g}^{\mathsf{T}},\forall g\in\mathcal{G}\},$
    $\mathcal{Y}_{\mathsf{forest-conn}}$ defined as
    $\mathcal{Y}_{\mathsf{forest-conn}}=\{\mathbf{Y}_{I}=\mathsf{blkdiag}(\mathbf{Y}_{I,1},\ldots,\mathbf{Y}_{I,G})~{}|$ (79)
[…]
    and $\mathcal{Y}_{\mathsf{band/stem-conn}}$ defined as
    $\mathcal{Y}_{\mathsf{band/stem-conn}}=\{\mathbf{Y}_{I}~{}|~{}\Re\{\mathbf{Y}_{I}\}=\mathbf{0}_{{M}\times{M}},$ (80)
[…]
@@
[26] Fully/Group-Connected Reflecting SISO Received Power Max.
@@
[51, 53] Propose closed-form global optimal solutions
[52] Fully-Connected
[40] Tree/Forest-Connected MISO, SIMO Propose a closed-form solution to achieve MISO optimum
[44] Dynamically Connected
@@
[25, 46] Non-Reciprocal Derive the optimal permuting strategy
[58] Multi-Sector Use non-reciprocal architectures to support multi-sector mode
[59] Fully-Connected Reflecting MIMO Capacity Max. Jointly design BD-RIS and transmitter covariance matrix
[60] Fully/Group-Connected
@@
[50] Non-Reciprocal Propose a Manifold-based solution with fast convergence
[61]
@@
[54] Fully/Group-Connected Reflecting Multi-User MISO Sum-Rate Max. Propose a two-stage design with low complexity
[64] Fully-Connected Propose a joint design based on unitary symmetric projection
[65] Fully/Group-Connected Propose a heuristic user scheduling scheme
[68, 69] Non-Reciprocal Propose learning-based beamforming design algorithms
[36] Hybrid Propose an iterative algorithm using Manifold theory
[43]
@@
[38] Group-Connected Multi-Sector Propose an iteratively closed-form solution
[66, 67]
@@
[70] Group-Connected
@@
[56] Fully/Group-Connected Reflecting Sum-Rate Max., Energy Effi. Max., Power Min.
@@ IV-C Channel Estimation
Channel estimation is crucial for BD-RIS-aided wireless systems since the performance benefits of BD-RIS are supported by proper optimization, beamforming design methods and performance analysis, all of which rely highly on accurate CSI. In D-RIS systems, there are generally two approaches to acquire instantaneous CSI, namely semi-passive channel estimation where RIS is mounted with RF chains for sensing signals, and passive channel estimation where RIS does not have the ability to sense signals [9, 16]. The semi-passive channel estimation approach results in separate base station-RIS and RIS-user channels with relatively low training overhead, both of which are independent of RIS architectures. Therefore, this approach used in D-RIS systems [71, 72, 73] is readily applicable to BD-RIS systems. However, the drawback of this approach is that the channel estimation performance relates tightly to the number of RF chains. Intuitively, the more the RF chains, the better the estimation performance. This indicates that a satisfactory estimation performance is achieved at the cost of significant power consumption. The passive channel estimation approach is an alternative approach which eases the requirement for RF chains [15]. Take the time-division duplex (TDD) system as an example, the passive channel estimation usually happens in uplink. Specifically, on the uplink, the user consecutively transmits pilot signals to the base station through BD-RIS and the BD-RIS keeps varying its response based on a pre-defined pattern. In this sense, it is in general not easy to obtain exactly the separate base station-RIS and RIS-user channels. Instead, the cascaded channel, which is the combination of base station-RIS and RIS-user channels, can be estimated. The passive channel estimation approach has been widely studied in D-RIS literature, starting by simple orthogonal RIS pattern design [74, 75, 76] with the aim to demonstrate the feasibility of this approach, followed by advanced pattern and protocol design with reduced training overhead [77, 78, 79]. Nevertheless, it is worth noting that the above passive channel estimation studies for D-RIS do not work for BD-RIS due to the following reasons.
@@
These two points will be explained in more details below, based on a specific communication system using the least-squares estimation.

### v2

@@ IV-B Survey on BD-RIS Optimization and Performance Analysis
Consider a general BD-RIS-aided multi-user MIMO system consisting of an $N$-antenna transmitter, an $M$-element BD-RIS, and $K$ multi-antenna users, each of which has $N_{k}$ antennas, $\forall k\in\mathcal{K}=\{1,\ldots,K\}$. Define the wireless channels from the transmitter to user $k$ as $\mathbf{H}_{{RT},k}\in\mathbb{C}^{N_{k}\times N}$, from the transmitter to BD-RIS as $\mathbf{H}_{{IT}}\in\mathbb{C}^{M\times N}$, and from the BD-RIS to user $k$ as $\mathbf{H}_{{RI},k}\in\mathbb{C}^{N_{k}\times M}$. According to [universal2024], assuming perfect matching, no mutual coupling, no specular reflection, and unilateral approximation at all devices, the overall channel $\mathbf{H}_{k}(\mathbf{\Theta})\in\mathbb{C}^{N_{k}\times N}$ from the transmitter to user $k$, which is a function of the scattering matrix $\mathbf{\Theta}$ of BD-RIS, writes as
@@
where $F(\mathbf{W},\{\mathbf{H}_{k}(\mathbf{\Theta})\}_{\forall k})$ denotes a general utility function varying according to performance metrics, $\mathcal{Y}$ denotes the constraint of $\mathbf{Y}_{I}$ varying according to architectures, and $\mathcal{W}$ denotes the constraint of the precoder matrix varying according to budget requirements. Note that (59) can be used to optimize all reciprocal architectures in Section III-D where $\mathbf{Y}_{I}$ is explicitly defined and constrained. Alternatively, the optimization problem can be formulated as the following scattering matrix based form
@@
where $\mathcal{T}$ denotes the constraints of $\mathbf{\Theta}$ and varying according to architectures. Note that (60) can be used to optimize single-, group-, fully-connected architectures in Section III-D and non-reciprocal architectures in Section III-E.
@@
For reciprocal BD-RIS, with the form of (60) and the constraint of dynamically connected architectures, [nerini2024static] has proposed an offline grouping strategy design adapting to the static CSI. For non-reciprocal BD-RIS, the following beamforming design and performance analysis studies have been conducted. With the form of (60) and a non-reciprocal BD-RIS architecture having non-diagonal scattering matrices being constrained by the set
@@
closed-form solutions can be derived. Focusing primarily on SISO systems with the utility function $F_{\mathsf{SISO}}(h(\mathbf{\Theta}))$, [li2022reconfigurable] has derived the performance upper-bound and the closed-form solution for BD-RIS. The beamforming design is later extended to MISO systems with the utility function $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$. With the form of (60), [li2024coordinated] has further proposed a coordinated non-reciprocal group-connected architecture with scattering matrices having more than $M$ non-zero entries, and derived the optimal grouping strategy which could maximize $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$. This architecture is further used to support the multi-sector mode and maximize $F_{\mathsf{MISO}}(\mathbf{w},\mathbf{h}(\mathbf{\Theta}))$ related to specific users [dong2024reconfigurable].
@@ IV-B.2 Capacity Maximization for MIMO
With the form of (60), the following beamforming design studies have been conducted. Focusing on reciprocal and fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$ ($G=1$)
@@
a common strategy is to iteratively design BD-RIS and the covariance matrix [santamaria2024mimo]. Specifically, the BD-RIS design can generally follow the idea of decoupling $\mathbf{\Theta}$ into $\mathbf{\Theta}=\mathbf{\Psi}^{\mathsf{H}}\mathbf{\Psi}$ with $\mathbf{\Psi}$ being unitary. For more specific LoS MIMO channels (both $\mathbf{H}_{RI}$ and $\mathbf{H}_{IT}$ have rank 1), the closed-form solution of $\mathbf{\Theta}$ can be obtained by exploiting the rank-1 property of channels $\mathbf{H}_{RI}$ and $\mathbf{H}_{IT}$ [santamaria2025rate]. Focusing on non-reciprocal BD-RIS architecture constrained by the set
@@
the problem has been solved by the Manifold method [zhao2024channel]. Matrix decomposition based solutions have also been proposed in closed form considering either near-field [bartoli2023spatial] or far-field [bjornson2024capacity] scenarios, which are proved to be globally optimal.
@@ IV-B.3 Sum-Rate Maximization for Multi-User MISO
This is again a multi-variable optimization problem, which can be solved either by separately designing BD-RIS and transmit precoder, or by block coordinate descent (BCD) methods [boyd2004convex].
For BD-RIS having reciprocal architectures, the following beamforming design and performance analysis studies have been conducted. With the form of (60) and group/fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$, a two-stage BD-RIS and precoder design has been proposed in [fang2023low] with low computational complexity. Specifically, the optimization of BD-RIS matrix is based on the symmetric unitary projection as detailed in Section IV-A. Following the idea of symmetric unitary projection, a joint precoder and BD-RIS design has been further proposed in [zhou2025joint]. With the form of (59) and group/fully-connected architectures constrained by $\mathcal{Y}_{\mathsf{group-conn}}$, a quasi-Newton based solution together with a heuristic user scheduling scheme has been proposed in [kim2023scattering]. With the form of (60) and multi-sector mode BD-RIS constrained by the set
@@
where $\varTheta_{l,1,n}$, $\forall l\in\mathcal{L}$, $\forall n=1,\ldots,\frac{M}{L}$ are extracted from $\mathbf{\Theta}$ according to Example 2 of Section III-F, an iteratively closed-form solution can be obtained by deriving the Karush-Kuhn-Tucker conditions [li2023beyond]. In addition, [samy2024enhancing] has derived the closed-form expression of the achievable sum-rate aided by multi-sector BD-RIS, with a further extension to other performance metrics, such as energy efficiency, error probability, outage probability [samy2024beyond].
For BD-RIS having non-reciprocal architectures, the following beamforming design studies have been conducted. For BD-RIS constrained by the set $\mathcal{T}_{\mathsf{non-recip}}$, [sobhi2024joint] and [loli2024meta] have proposed learning-based methods which support larger-dimensional optimizations. For BD-RIS with hybrid mode and proper non-reciprocal architectures, the constraint $\mathbf{\Theta}_{\mathsf{r}}=\mathbf{\Theta}_{\mathsf{r}}^{\mathsf{T}}$ from Example 1 of Section III-F can be dropped such that $\mathbf{\Theta}$ is constrained by the set
@@
where $\mathbf{\Theta}_{\mathsf{r},g}$ and $\mathbf{\Theta}_{\mathsf{t},g}$, $\forall g\in\mathcal{G}$ are extracted from $\mathbf{\Theta}$ according to Example 1 of Section III-F. It should be noted here that the constraint of hybrid BD-RIS essentially constructs a Stiefel Manifold [lee2018introduction]. Therefore, an iterative searching-based solution can be obtained by using the Manifold theory [li2022beyond]. For BD-RIS with hybrid mode and dynamically connected architectures adapting to the instantaneous CSI, [li2023dynamic] has proposed a heuristic grouping strategy design.
@@ IV-B.4 Energy Efficiency Maximization and Power Minimization for Multi-User MISO
For BD-RIS having reciprocal architectures, the following beamforming design and performance analysis studies have been conducted. With the form of (60) and group/fully-connected architectures constrained by $\mathcal{T}_{\mathsf{group-conn}}$, [zhou2023optimizing] has proposed an iterative framework by decoupling the unitary and symmetric constraints of $\mathbf{\Theta}$ with a newly introduced auxiliary variable $\mathbf{\Phi}=\mathbf{\Theta}$, such that $\mathbf{\Theta}$ is only subject to the unitary constraint and $\mathbf{\Phi}$ is only subject to the symmetric constraint, or vise versa. The proposed framework is applicable to sum-rate maximization, energy-efficiency maximization, and power minimization problems. With the form of (60) and STARS, which is essentially a special case of multi-sector BD-RIS constrained by $\mathcal{T}_{\mathsf{multi-sec}}$ when $L=2$ and also a special case of hybrid BD-RIS constrained by $\mathcal{T}_{\mathsf{hyb}}$ when $\bar{M}=2$, [farhadi2025meta] has proposed a meta-learning approach for joint optimization of BD-RIS, resource allocation, and antenna selection. With the form of (59), [wu2024optimization] has recently proposed a universal framework applicable to arbitrary constraints of BD-RIS admittance matrix, the aforementioned three utility functions, and sum-rate maximization for more general multi-user MIMO scenarios. The main idea is to introduce auxiliary variables $\mathbf{u}_{k}=\mathbf{h}_{RI,k}\mathbf{\Theta}$, $\forall k\in\mathcal{K}$, such that the matrix inverse in (59) can be eliminated by transferring it to bilinear constrains to facilitate the optimization. The applicable BD-RIS constraints include but are not limited to those for group/fully-, forest/tree-, and band/stem-connected architectures in Section III-D.
@@
[shen2021] Fully/Group-Connected Reflecting SISO Received Power Max.
@@
[nerini2023closed, santamaria2023snr] Propose closed-form global optimal solutions
[sun2024new] Fully-Connected
[nerini2024beyond] Tree/Forest-Connected MISO, SIMO Propose a closed-form solution to achieve MISO optimum
[nerini2024static] Dynamically Connected
@@
[li2023reconfigurable, li2024coordinated] Non-Reciprocal Derive the optimal permuting strategy
[dong2024reconfigurable] Multi-Sector Use non-reciprocal architectures to support multi-sector mode
[santamaria2024mimo] Fully-Connected Reflecting MIMO Capacity Max. Jointly design BD-RIS and transmitter covariance matrix
[santamaria2025rate] Fully/Group-Connected
@@
[zhao2024channel] Non-Reciprocal Propose a Manifold-based solution with fast convergence
[bartoli2023spatial]
@@
[fang2023low] Fully/Group-Connected Reflecting Multi-User MISO Sum-Rate Max. Propose a two-stage design with low complexity
[zhou2025joint] Fully-Connected Propose a joint design based on unitary symmetric projection
[kim2023scattering] Fully/Group-Connected Propose a heuristic user scheduling scheme
[sobhi2024joint, loli2024meta] Non-Reciprocal Propose learning-based beamforming design algorithms
[li2022beyond] Hybrid Propose an iterative algorithm using Manifold theory
[li2023dynamic]
@@
[li2023beyond] Group-Connected Multi-Sector Propose an iteratively closed-form solution
[samy2024enhancing, samy2024beyond]
@@
[farhadi2025meta] Group-Connected
@@
[zhou2023optimizing] Fully/Group-Connected Reflecting Sum-Rate Max., Energy Effi. Max., Power Min.
@@ IV-C Channel Estimation
Channel estimation is crucial for BD-RIS-aided wireless systems since the performance benefits of BD-RIS are supported by proper optimization, beamforming design methods and performance analysis, all of which rely highly on accurate CSI. In D-RIS systems, there are generally two approaches to acquire instantaneous CSI, namely semi-passive channel estimation where RIS is mounted with RF chains for sensing signals, and passive channel estimation where RIS does not have the ability to sense signals [zheng2022survey]. The semi-passive channel estimation approach results in separate base station-RIS and RIS-user channels with relatively low training overhead, both of which are independent of RIS architectures. Therefore, this approach used in D-RIS systems [hu2021semi, alexandropoulos2020hardware, taha2019deep] is readily applicable to BD-RIS systems. However, the drawback of this approach is that the channel estimation performance relates tightly to the number of RF chains. Intuitively, the more the RF chains, the better the estimation performance. This indicates that a satisfactory estimation performance is achieved at the cost of significant power consumption. The passive channel estimation approach is an alternative approach which eases the requirement for RF chains [swindlehurst2022channel]. Take the time-division duplex (TDD) system as an example, the passive channel estimation usually happens in uplink. Specifically, on the uplink, the user consecutively transmits pilot signals to the base station through BD-RIS and the BD-RIS k.pdf varying its response based on a pre-defined pattern. In this sense, it is in general not easy to obtain exactly the separate base station-RIS and RIS-user channels. Instead, the cascaded channel, which is the combination of base station-RIS and RIS-user channels, can be estimated. The passive channel estimation approach has been widely studied in D-RIS literature, starting by simple orthogonal RIS pattern design [yang2020intelligent, zheng2019intelligent, you2020channel] with the aim to demonstrate the feasibility of this approach, followed by advanced pattern and protocol design with reduced training overhead [wang2020channel, guan2021anchor, zhou2022channel]. Nevertheless, it is worth noting that the above passive channel estimation studies for D-RIS do not work for BD-RIS due to the following reasons.
@@
These two points will be explained in more details, based on a specific least-squares estimation.